10) The market portfolio has an expected return of 10% and standard deviation of returns of 15%. The riskless interest rate is 4%. What is the maximum standard deviation an investor should accept in order to earn an expected return of 16%?

a) 15% b) 20% c) 25% d) 30% e) None of the above
Answer #1

Expected Seturn of market poortfolio (Rm) - 10:7. Standard deviation of manket com] = 15.1. risk less interest rate [Rf] = 4%© = 16 = 4 + 25 (10-4). TETE 15 =) 2) 16 = 4+ (1.6 16 = 13.99 so optiow © worong. 9934 08 as to go @ =) 16 = 4 + 30 (10-4) +